OpenStudy (anonymous):

X-4/2x-6+3/x^2-9 Which one is the simplified form? A. X+2/2(x+3) B. x-2/2(x-3) C. X-2/2(x+3) D. X+2/2(x-3)

OpenStudy (anonymous):

is it that?? \[\frac{ x-4 }{ 2x-6 }+\frac{ 3 }{ x ^{2}-9 }\]

OpenStudy (anonymous):

\[\frac{ (x+3)(x-4) }{ (x+3)(2x-6) }+\frac{ 2*3 }{ 2*(x ^{2}-9) }\] \[\frac{ x ^{2}-x-6 }{ 2*(x-3)(x+3) }\] (x+2)/2*(x+3) answer is A.